# ALIFE2024 template

### v1

@@ Abstract
A major challenge when describing the origin of life is to explain how instructional information control systems emerge naturally and spontaneously from mere molecular dynamics. So far, no one has clarified how information control emerged ab initio and how primitive control mechanisms in life might have evolved, becoming increasingly refined. Based on recent experimental results showing that chemical computation does not require the presence of life-related chemistry (Dueñas-Díez and Pérez-Mercader,, 2019), we elucidate the origin and early evolution of information handling by chemical automata, from information processing (computation) to information storage (memory) and information transmission (communication). In contrast to other theories that assume the existence of initial complex structures, our narrative starts from trivial self-replicators whose interaction leads to the arising of more powerful molecular machines. By describing precisely the primordial transitions in chemistry-based computation, our metaphor is capable of explaining the above-mentioned gaps and can be translated to other models of computation, which allow us to explore biological phenomena at multiple spatial and temporal scales. At the end of our manuscript, we propose some ways to extend our ideas, including experimental validation of our theory (both in vitro and in silico).
@@
As was explained in Walker and Davies, (2013), neither genetic-first nor metabolism-first approaches can describe completely the emergence of living organisms, as both viewpoints neglect the active and distributed nature of biological information. In this way, the real challenge when describing life’s origin is to explain how instructional information control systems emerge naturally and spontaneously from simple molecular dynamics. Thus, we explore an information-first approach to study the origins and evolution of life. Recently, it was shown that chemical reactions can be thought of as molecular recognition machines, which do not need biochemistry to mimic the behavior of automata belonging to different classes (Dueñas-Díez and Pérez-Mercader,, 2019).
@@
The experiments carried out by Dueñas-Díez and Pérez-Mercader, (2019) (of which the second author will be one of the keynote speakers at ALIFE 2024) consisted of implementing strictly native chemical realizations of automata in their classical hierarchy. To enact a Finite-State Automaton (FSA), they used a simple bimolecular reaction in an aqueous medium. To render a one-stack Push-Down Automaton (PDA) they employed a pH-reaction network. To bring about a Turing machine in the Limited Bounded Automaton (LBA) subclass, they used the Belousov-Zhabotinsky (BZ) reaction network. As the authors have pointed out in their conclusions, these experimental results enlighten a path for the sequential/hierarchical origin of information handling in nature and the transition from non-biochemical to biochemical and extant life-related computation.
@@
There are several theoretical frameworks which describe life as information (Chaitin,, 2012; Farnsworth et al.,, 2013; Marletto,, 2015), but none of them has explained how information handling emerged from chemical interactions. By information handling, we mean information processing (computation), information storage (memory), and information transmission (communication). In this work, we describe precisely the primordial transitions in chemistry-based computation, from the appearance of simple trivial replicators (FSAs) up to the evolution of chemical Turing machines (LBAs) and, eventually, to their combination into more powerful molecular computers.
@@
For us, FSAs demarcate the origin of information processing (computation), PDAs define the appearance of information storage (memory) and LBAs bound the emergence of information transmission (communication). This correspondence is depicted in Fig. 1. By elucidating the origin and early evolution of information handling by chemical automata, in the next three sections we explain how information control emerged ab initio and how primitive control mechanisms in life might have evolved. We do not assume the existence of initial complex structures and our narrative can be translated to other models of computation, which in future work would allow us to explore biological phenomena at multiple spatial and temporal scales.
@@
Figure 1: Relationship between classes of automata and the origin of the different types of information handling. The large circles represent the three forms of information handling: computation (information processing), memory (information storage), and communication (information transmission). FSAs are only capable of processing information. PDAs also have memory capacity. LBAs can additionally transmit information. For each class of automata we checklist some important properties that they exhibit given our narrative.
@@
We can imagine our planet as a big parallel, quasi-universal Turing machine capable of simulating a huge variety of functions (programs). The term “quasi-universal” is due to the fact that “real” universal Turing machines are pure abstractions which need an infinite memory tape. Of course, natural resources are finite, and this is a first limitation for life to emerge. Under this perspective, the first organisms on the planet could be described as some kind of automata which were running in these primitive environments. As a result of the geophysical and geochemical constraints, which we could call the primeval ecosystem language (Pattee,, 1969), elementary bimolecular reactions emerged. These chemical reactions are capable of recognizing regular languages, a task that does not involve counting or memory (Dueñas-Díez and Pérez-Mercader,, 2019).
These trivial replicators, as Walker and Davies, (2013) call them, are simple enough to have been naturally occurring in pre-biological environments and kick off the origin of information processing. These structures do not require a memory nor a “vehicle” (in the von Neumann sense), so we can imagine the architecture of these earliest chemical machines as being conformed by a single FSA, capable of constructing itself using natural resources and a simple description. Just as $L$-systems, it is clear that the algorithmic information of these simple architectures is one that may be specified by an algorithm containing far fewer bits than the structure it describes.
@@
In this primitive environment, it is possible to achieve self-replication (a process that differs from self-reproduction (Marletto,, 2015)) by simple bimolecular reactions (Duim and Otto,, 2017). Since it involves the unassisted formation of a covalent bond between the two reactants, the reactions will be relatively slow. Nevertheless, the autocatalytic cycle will play an increasingly dominant role (Kauffman,, 1986) if a sufficiently large amount of the template molecules is formed (Duim and Otto,, 2017). Although it has been speculated that the simple interaction between self-replicators and the environment leads to open-ended evolution, for now this is not a prerequisite.
@@
So far, we have a world in which molecular FSAs coexist. These are capable of computation and self-replication but have no vehicles or memory. The diversity of molecules in the environment allows them to interact and, if possible, bind with each other. This generates systems of multiple reactions that are interconnected through common intermediate chemical species. Under this panorama, some chemical intermediates are acting as the product of one reaction and as reactant of another one. This dynamical process is essentially equivalent to the push/pop duality in one-stack PDAs (Dueñas-Díez and Pérez-Mercader,, 2019). This transition from FSAs to one-stack PDAs demarcates the appearance of analog structures, in which delocalized information storage (memory) allows for a higher evolutionary robustness in the face of changing environmental conditions (Walker and Davies,, 2013).
@@
From a von Neumann mindset, the one-stack PDA is now some kind of primitive constructor. At this computational level, self-replication is the result of nature trying to solve a equivalent PDA-problem to the Halting problem (Sayama,, 2008). In this case, the ecosystem, a quasi-universal computer compartment, tries to predict which PDA will be the longest-running machine. This is equivalent to computing the largest number of steps taken by an $n$-state, $m$-symbol machine started on an initially non-blank tape before halting. This is of course an undecidable problem, and at the level of the PDA, when this attempts to execute the instruction of being the longest-running machine, it produces a nested loop of computation which in terms of construction translates as self-replication. This process is summarized in Fig. 2.
@@
Figure 2: Explanation of PDA and LBA self-replication. First, the environment (described as a quasi-universal computer compartment) tries to predict which replicator will be the longest-running machine. Next, the automata attempt to execute the instruction of being the longest-running machine. This results in a nested loop of computation which in terms of construction translates as self-replication (Sayama,, 2008). Once the automata self-replicate, the cycle repeats. Undecidability explicitly leads us to have open-ended evolution. (Rocha,, 1998; Hernández-Orozco et al.,, 2018).
@@
The undecidability behind self-replication lead us explicitly to open-ended evolution (Rocha,, 1998; Hernández-Orozco et al.,, 2018). In fact, the delocalized memory allows the PDA to store a more complex set of instructions, including the components of the chemical network conforming the PDA itself. Thus, in this evolutionary phase the system is capable of developing extra functions, which gives to the PDA the ability of self-preserving behavior and propagate mutations to its offspring. In addition, observe that the transition between FSAs and one-stack PDAs is compatible with ideas related to the appearance composomes (Segré and Lancet,, 2000; Hunding et al.,, 2006) and their maintenance (Markovitch and Lancet,, 2012).
@@
At this stage we can talk about ante-organisms, life-like systems that are capable of self-preserving behaviour (Egbert et al.,, 2023). The minimal requirement for natural selection is that each kind of replicator produce at least one viable offspring, on average, per lifetime (Marletto,, 2015). Therefore, we can claim that so far, our ante-organisms are capable of primitive evolution (some kind of natural selection with limited heredity). Of course, it is well-known that analog-only systems are not as versatile or as robust as analog systems with digital information control (Walker and Davies,, 2013), and as such, may probably have a limited evolutionary capacity (Vasas et al.,, 2010).
@@
The next level of computational complexity are Turing machines, a class of automaton equivalent to a two or more stacks PDA (Hopcroft et al.,, 2001). Following our train of thought, there are two main, non-exclusive ways for obtaining an LBA. The first one is through the complexification of one initial PDA. Since at this stage the entities are able to propagate mutations across their offspring, it is possible that they can become more and more complex, in a way that they can develop two or more push/pop dualities. The second way is essentially through the coupling of two or more PDAs. Our model does not exclude the coexistence of multiple chemical automata with different degrees of complexity. Given an environment with finite resources promoting undecidability, these different organisms can cooperate to persist.
@@
Because chemical automata need communication in some way to cooperate, this era launches information transmission. At this evolutionary phase, we can talk about some kind of compartmentalization, in the sense that now the chemical constructors can be made of at least two other chemical subconstructors, increasing their functionality and robustness. In this period, we might have observed the emergence of components capable of copying non-digital patterns and performing part of the manipulation of those rules. Notice that at this stage the abstract and non-physical systemic entities effectively become causal agents capable of manipulating more and more of its material substrate (Auletta et al.,, 2008). Following the mindset of (Walker and Davies,, 2013), we are having a transition from trivial to non-trivial replicators, where the latter requires an algorithm, or instruction set, of complexity comparable to the system it describes (or creates).
@@
Since undecidability is plausible in the computational classes of PDA and LBA (Rado,, 1962), the entities can still self-replicate by the mechanism previously described (see Fig. 2). However, they are not yet capable of genetic-like Darwinian evolution, so we still have ante-organisms (Egbert et al.,, 2023). In this world of FSAs, PDAs, and LBAs interacting in a dynamical environment, the appearance of digital messengers is not totally unlikely since these structures should have had enough time to emerge as the product of some synergistic process, in much the same way that our formal rules of syntax and dictionaries were condensed from the functional usage of primitive symbols in a complex environment (Pattee,, 1969). Along these lines, digitization would have arisen as a natural outcome of the processes in reaction networks that had once been primarily analog. This fact helped automata to communicate more efficiently among them and at the same time with their offspring. Certainly, this process could have taken a long time to happen, such that digital life forms may be the only systems that survive on the long-run and are thus the only remaining product of the processes that led to life (Walker and Davies,, 2013).

### v2

@@ Abstract
A major challenge when describing the origin of life is to explain “how instructional information control systems emerge naturally and spontaneously from mere molecular dynamics” (Walker and Davies,, 2013). So far, no one has clarified how information control emerged ab initio and how primitive control mechanisms in life might have evolved, becoming increasingly refined. Based on recent experimental results showing that chemical computation does not require the presence of life-related chemistry (Dueñas-Díez and Pérez-Mercader,, 2019), we elucidate the origin and early evolution of information handling by chemical automata, from information processing (computation) to information storage (memory) and information transmission (communication). In contrast to other theories that assume the existence of initial complex structures, our representation starts from trivial self-replicators whose interaction leads to the arising of more powerful molecular machines. By describing precisely the primordial transitions in chemistry-based computation, our framework is capable of explaining the above-mentioned gaps and can be translated to other models of computation, which allow us to explore biological phenomena at multiple spatial and temporal scales. At the end of our manuscript, we propose some ways to extend our ideas, including experimental validation of our theory (both in vitro and in silico).
Introduction
Leaving aside the philosophical question about the origin of information or the chicken-and-egg problem between information and matter/energy, it is possible to have an even deeper understanding of phenomena such as life or cognition if we assume that information is anything that an entity can sense, perceive, or observe. (Gershenson,, 2012). Thus, following the line of thought proposed by Pattee, (1969), we can establish the primordial ecosystem as the first information-containing architecture in our planet, whose primitive geochemical cycles gave way to a well-defined syntax—based on chemistry— spontaneously without the need for genetic instructions or initial metabolic control. This gave way to molecular Darwinism, a phenomenon that has been well documented for several decades (Küppers,, 1990) and whose formalization was refined the last decade by Marletto, (2015).
@@ Introduction
As was explained in Walker and Davies, (2013), neither genetic-first nor metabolism-first approaches can describe completely the emergence of living organisms, as both viewpoints “neglect the active (algorithmic or instructional) and distributed nature of biological information” (Walker and Davies,, 2013). In this way, “the real challenge of life’s origin is thus to explain how instructional information control systems emerge naturally and spontaneously from mere molecular dynamics” (Walker and Davies,, 2013). Thus, we explore an information-first approach to study the origins and evolution of life. Recently, it was shown that chemical reactions can be thought of as molecular recognition machines, which do not need biochemistry to mimic the behavior of recognizing automata belonging to different classes in Chomsky’s hierarchy (Dueñas-Díez and Pérez-Mercader,, 2019).
@@
The experiments carried out by Dueñas-Díez and Pérez-Mercader, (2019) consisted of implementing strictly “one-pot-reactor” chemical realizations of automata for the recognition of well-known languages at different levels in the Chomsky hierarchy. To enact a Finite-State Automaton (FSA), they used a simple bimolecular reaction in an aqueous medium. Such reaction was capable of recognizing a regular language. To render a one-stack Push-Down Automaton (PDA) they employed a pH-reaction network, being capable of recognizing a context-free language. To bring about a Turing machine in the Linear Bounded Automaton (LBA) subclass, they used the Belousov-Zhabotinsky (BZ) reaction network. In this case, the BZ reaction was able to successfully recognize a context-sensitive language. As the authors have pointed out in their conclusions, these experimental results enlighten “a path for the sequential/hierarchical origin of information handling in nature and the transition from non-biochemical to biochemical and extant life-related computation” (Dueñas-Díez and Pérez-Mercader,, 2019).
@@
There are several theoretical frameworks which describe life as information (Chaitin,, 2012; Farnsworth et al.,, 2013; Marletto,, 2015), but none of them has explained how information handling emerged from chemical interactions. By information handling, we mean information processing (computation), information storage (memory), and information transmission (communication). As will be discussed below, such information handling also includes the syntactic, semantic and pragmatic flavors of information. In this work, we describe precisely the primordial transitions in chemistry-based computation, from the appearance of trivial replicators (simple bimolecular reactions, equivalent to FSAs) in the environment up to the evolution of approximated chemical Turing machines (complex chemical networks, represented here by LBAs) and, eventually, to their combination into more powerful molecular computers, capable of performing Bayesian inference (Poole et al.,, 2017) and associative learning (Bartlett and Louapre,, 2022), which could have led to the development of autonomy (Kirchhoff et al.,, 2018) and high-level cognition (Levin,, 2019).
@@
For us, the presence of FSAs in the environment demarcates the origin of information processing (computation), PDAs define the appearance of information storage (memory) by the environment and LBAs bound the emergence of information transmission (communication) through different entities in the environment. This correspondence is depicted in Fig. 1. By elucidating the origin and early evolution of information handling by chemical automata, in the next three sections we explain how information control emerged ab initio and how primitive control mechanisms in life might have evolved. We do not assume the existence of initial complex structures and our narrative can be translated to other models of computation, which in future work would allow us to explore biological phenomena at multiple spatial and temporal scales.
@@
Figure 1: Relationship between classes of automata and the origin of the different types of information handling. The large circles represent the three forms of information handling: computation (information processing), memory (information storage), and communication (information transmission). FSAs allow the environment to process information. PDAs permit a primitive form of memory capacity by the environment. LBAs empower communication between different entities in the environment. For each class of automata we checklist some important properties that they grant to the ecosystem given our framework.
@@
As has been suggested by Chaitin, (2012), think of our planet as a big parallel, quasi-universal Turing machine capable of simulating a huge variety of functions (programs) can give us novel insights about evolution. Here, the term “quasi-universal” is due to the fact that “real” universal Turing machines are pure abstractions which need an infinite memory tape. Natural resources are finite, and this is a first limitation for life to emerge. Extrapolating the ideas suggested in Hernández-Orozco et al., (2018) and inspired by the evolutionary model proposed in von Neumann, (1966), the first organisms on the planet could be described as some kind of primitive automata which were assembled and executed by these primeval habitats. As noted above, the primordial geophysical and geochemical constraints generated the primeval ecosystem language (Pattee,, 1969), which was interpreted by such primitive entities. The cyclic nature of such processes allows for a robust flow of information—abolishing the fragility or deterioration of it—, resulting in the emergence of elementary autocatalytic bimolecular reactions. These chemical reactions are capable of recognizing regular languages, a task that does not involve counting or memory (Dueñas-Díez and Pérez-Mercader,, 2019).
@@
By the good regulator theorem (Conant and Ross Ashby,, 1970), some of the elementary bimolecular reactions in that primeval environment must have a cyclic nature similar to that of geologic phenomena. Thus, it is possible to achieve self-replication by simple bimolecular reactions. Since it involves the unassisted formation of a covalent bond between the two reactants, the reactions will be relatively slow (Duim and Otto,, 2017). Nevertheless, the autocatalytic cycle will play an increasingly dominant role (Kauffman,, 1986) if a sufficiently large amount of the template molecules is formed (Duim and Otto,, 2017), which in part could also promote self-replication. Self-replication differs from self-reproduction in the existence of a copier $C$ acting on raw material (Szathmáry,, 1995). As has been well argued by Marletto, (2015): “for looser replicators, such as crystals, or short RNA strands and autocatalytic cycles involved in the origin of life, the copier $C$ is null, i.e. implicit in the laws of physics”. Although it has been speculated that the simple interaction between self-replicators and the environment leads to open-ended evolution, for us this is not a prerequisite.
These trivial replicators, as Walker and Davies, (2013) call them, are simple enough to have been naturally occurring in pre-biological environments and kick off the origin of information processing. These structures do not require a memory nor a “vehicle”—in the von Neumann sense (von Neumann,, 1966)—, so we can imagine the architecture of these earliest chemical machines as being represented by autocatalytic bimolecular chemical reactions, which under our framework are single FSAs (Dueñas-Díez and Pérez-Mercader,, 2019), capable of constructing themselves using natural resources and a simple description encoded in the environment. Just as $L$-systems, it is clear that the algorithmic information of these simple architectures is one that may be specified by an algorithm containing far fewer bits than the structure it describes.
At this point, it is important to note that the different dimensions of the concept of information (syntactic, semantic and pragmatic) cannot be separated from each other (Küppers,, 1990). Although the existence of semantic (or pragmatic) information in such a primordial state of life may seem far-fetched, it must first be remembered that meaning is a multi-scale phenomenon (Fields and Levin,, 2020). According to Küppers, (1990): “the protosemantics of biological information is clearly defined by the ability of the biological information carrier to reproduce as quickly as possible while maintaining high accuracy and stability”, which is fully in line with our hypothesis. In fact, it has been shown that it is possible to approximate Bayesian inference using simple bimolecular reactions (Poole et al.,, 2017), which could have promoted the decoding of symbol-to-meaning information and subsequently developed semantic closure (Pattee,, 1995).
@@
So far, we have a world in which bimolecular reactions (represented here by FSAs) coexist. These are capable of computation and self-replication but have no vehicles or memory. The diversity of molecules in the environment allows them to interact and, if possible, bind with each other. This generates systems of multiple reactions that are interconnected through common intermediate chemical species. Under this panorama, some chemical intermediates are acting as the product of one reaction and as reactant of another one, forming networks of even more complex chemical reactions. This dynamical process is identical to the premise of autocatalysis and is essentially equivalent to the push/pop duality in one-stack PDAs described in (Dueñas-Díez and Pérez-Mercader,, 2019). The transition from FSAs (simple bimolecular reactions) to one-stack PDAs (networks of chemical reactions) demarcates the appearance of a primitive information storage (memory) by the environment, because it allows for the recognition of more complex sequences of information (Hopcroft et al.,, 2001) and thus it empowers a higher evolutionary robustness of such information in the face of changing environmental conditions (Walker and Davies,, 2013).
From a von Neumann mindset, the one-stack PDA is now some kind of primitive constructor, because we have a chemical network that might be able to reproduce itself including the components that conform it. At this computational level, self-replication can be seen as the result of nature trying to solve a equivalent PDA-problem to the Halting problem (Sayama,, 2008). Following our computational train of thought, the ecosystem, a quasi-universal computer compartment, “is constantly approximating solutions to the adaptation problem by means of survivability” (Hernández-Orozco et al.,, 2018). In this way, nature pursuits of the busy beaver values (longest-running automata) in the environment, which is identical to the search for the longest-surviving organisms.
@@
At this particular evolutionary stage, nature is looking for which PDA will be the longest running machine, which is interchangeable to find the most robust primitive information storage for the environment. This is equivalent to computing the largest number of steps taken by an $n$-state, $m$-symbol machine started on an initially non-blank tape before halting. This is of course an undecidable problem comparable to a busy beaver game (Rado,, 1962), and at the level of the PDA, when this attempts to execute the instruction of being the longest-running machine, it produces a nested loop of computation which in terms of construction translates as self-replication (Sayama,, 2008). This process is summarized in Fig. 2.
@@
Figure 2: Explanation of PDA and LBA self-replication. First, the environment (described as a quasi-universal computer compartment) approximates solutions to the adaptation problem by means of survivability, which guides the search for the longest-surviving organisms. Next, the automata (a chemical network) attempt to execute the instruction of being the longest-running machine. This results in a nested loop of computation which in terms of construction translates as self-replication (Sayama,, 2008). Once the automata self-replicate, the cycle repeats. Undecidability explicitly leads us to have open-ended evolution. (Hernández-Orozco et al.,, 2018).
@@
The undecidability behind self-replication lead us explicitly to open-ended evolution as a result of evolution (Pattee and Sayama,, 2019), and as Hernández-Orozco et al., (2018) has pointed out, systems that exhibit (strong) open-ended evolution must be undecidable. The PDA allows the ecosystem to store a more complex set of instructions—including the components of the chemical network that makes up the PDA itself—in a delocalized way. Thus, in this evolutionary phase the system is capable of developing extra functions, which gives to the PDA the ability of self-preserving behavior and propagate mutations to its offspring. This is consistent with the fact that a simple catalytic reaction network may be able of encoding correlations in environmental variables, evolving some kind of associative learning (Bartlett and Louapre,, 2022). Following the law of requisite variety (Ashby,, 1956), these chemical networks evolve concurrently to the environment in which they were developed, producing more semantic information which translates as a primitive degree of agency (Kolchinsky and Wolpert,, 2018). In addition, observe that the transition between FSAs and one-stack PDAs is compatible with ideas related to the appearance composomes (Segré and Lancet,, 2000; Hunding et al.,, 2006) and their maintenance (Markovitch and Lancet,, 2012).
@@
At this stage we can talk about ante-organisms, life-like systems that are capable of self-preserving behaviour (Egbert et al.,, 2023). Up to this point we have a resource-limited ecosystem inhabited by chemical networks that compete, self-replicate and mutate (through small variations in their composition) in such a way that each kind of replicator produce at least one viable offspring, on average, per lifetime. Therefore, we can claim that so far, our ante-organisms are capable of primitive evolution (Marletto,, 2015), which is a kind of natural selection with limited inheritance caused by low robustness and accuracy in replicating the information stored in the chemical network by the environment. However, according to Walker and Davies, (2013): “analogue-only systems are not as versatile or as robust as analogue systems with digital information control, and as such, may probably have a limited evolutionary capacity (Vasas et al.,, 2010)”.
@@
The next level of computational complexity are Turing machines, a class of automaton equivalent to a two or more stacks PDA (Hopcroft et al.,, 2001). Following our train of thought, there are two main, non-exclusive ways for obtaining an LBA. The first one is through the complexification of one initial PDA. Since at this stage the entities are able to propagate mutations across their offspring (represented by variations in the structures of the chemical networks), it is possible that they can become more and more complex, in a way that they can develop two or more push/pop dualities. The second way is essentially through the coupling of two or more PDAs. Our model does not exclude the coexistence of multiple chemical automata with different degrees of complexity. Given an environment with finite resources promoting survivability, these different organisms can cooperate to persist. In both cases we obtain even more complex chemical networks that allow the identification of more abstract sequences of information, resulting in an even more sophisticated information storage system on the part of the environment.
@@
Because chemical automata need communication in some way to cooperate, this era launches information transmission. Given a set of selfish agents minimizing their stress (surprise) and competing for information, it is possible to develop cooperative homeostatic loops which improve the predictive capabilities of the entities and gives rise to functional relationships (Levin,, 2019; Heins et al.,, 2024). Such infotaxis may have resulted in some kind of compartmentalization, in the sense that now the chemical constructors can be made of at least two chemical networks, increasing their functionality and robustness. Notice that at this stage the abstract and non-physical systemic entities effectively become causal agents capable of manipulating more and more of its material substrate (Auletta et al.,, 2008). Following the mindset of (Walker and Davies,, 2013), we are having a transition from trivial to non-trivial replicators, where the latter requires an algorithm, or instruction set, of complexity comparable to the system it describes (or creates). In the same way, we can imagine this metamorphosis from reactive to representational systems (Zenil et al.,, 2012).
@@
Since undecidability is plausible in the computational classes of PDA and LBA (Rado,, 1962), the entities can still self-replicate by the mechanism previously described (see Fig. 2). However, they are not yet capable of genetic-like Darwinian evolution, so we still have ante-organisms (Egbert et al.,, 2023). In this world of FSAs, PDAs, and LBAs interacting in a dynamical environment, the appearance of digital messengers is not totally unlikely since these structures should have had enough time to emerge as the product of some synergistic process, in much the same way that our formal rules of syntax and dictionaries were condensed from the functional usage of primitive symbols in a complex environment (Pattee,, 1969). Along these lines, digitization would have arisen as a natural outcome of the processes in reaction networks that had once been primarily analog. This fact helped automata to communicate more efficiently among them and at the same time with their offspring. Certainly, this process could have taken a long time to happen, such that digital life forms may be “the only systems that survive in the long-run and are thus the only remaining product of the processes that led to life” (Walker and Davies,, 2013).